An object's position is given by the vector s⃗ (t)=2cos(ωt)i^+2sin(ωt)j^whereω=1radianpersecond.(there are hats on top of i and k in equation) How would you describe this object's motion? Hint: if you are stuck, compute and graph its position in the Cartesian plane at a variety of values of t (say, for integers 0-6).
Answer #1

Here ,

the position of the vector is s(t)

s(t) = 2 * cos(w *t) i + j * sin(w*t)

x = 2 * cos(w*t)

y= 2 *sin(w*t)

squaring and adding

x^2 + y^2 = 4 * cos^2(w*t) + 4 * sin^2(w * t)

x^2 + y^2 = 4

the object will move in a circle with centre at origin and radius 2